In a random sample of 13 residents of the state of Florida, the mean waste recycled per person per day was 2.9 pounds with a standard deviation of 0.94 pounds. Determine the 98% confidence interval for the mean waste recycled per person per day for the population of Florida. Assume the population is approximately normal.

Step 1 of 2 :  

Find the critical value that should be used in constructing the confidence interval. Round your answer to three decimal places.
Answer #1

Solution :

sample size = n = 13

Degrees of freedom = df = n - 1 = 12

t /2,df = 2.681

Margin of error = E = t/2,df * (s /n)

= 2.681 * (0.94 / 13)

Margin of error = E = 0.70

The 98% confidence interval estimate of the population mean is,

- E < < + E

2.9 - 0.70 < < 2.9 + 0.70

2.20 < < 3.60

(2.20 , 3.60)